12. Phosphorus pentachloride can be produced by the reaction of phosphorus trichloride and chlorine gas. What mass of phospho Please provide all steps and a detailed explanation
Answer #1

PCl5(g)----------------->PCl3(g)   + Cl2(g)

1mole of PCl5 decomposes to gives 1 mole of PCl3

208.24g of PCl5 decomposes to gives 137.33g of PCl3

127g of PCl5 deomposes to gives = 137.33*127/208.24   =83.75g of PCl3

percent yiled = actual yield*100/theoretical yield

   84.8              = actual yield*100/83.75

actual yield    = 84.8*83.75/100   = 71.02g of PCl3